OpenStudy (steve816):

Please help me how to set this up. I don't know how. A 1-kg block of wood is dropped from a height 2.26 meters above a resting 0.05-kg ball. Shortly thereafter, the ball is shot horizontally at a velocity of 2.91 m/s. The wood is 9.94 meters away from the ball in the horizontal direction. Determine the distance (in meters) the wood falls from its original position until impact? Ignore air resistance.

OpenStudy (518nad):

2.26m?

OpenStudy (steve816):

Nope, was wrong :/

OpenStudy (518nad):

i was assuming the ball was on teh ground, but i guess not

OpenStudy (518nad):

9.94/2.91 secs the time wood was being dropped for

OpenStudy (518nad):

a=9.81m/s v=9.81 * t d=9.81/2*t^2 d=9.81/2*(9.94/2.91)^2

OpenStudy (flcon04):

do you still need help?

OpenStudy (steve816):

Yes @flcon04